\begin{thebibliography}{[CLOT]a}
\bibliographystyle{amsalpha}


\bibitem[Ba93]{Bab1} Babenko, I.: Asymptotic invariants of smooth
manifolds.  {\em Russian Acad. Sci. Izv. Math.} \textbf{41} (1993),
1--38.


\bibitem[BKSW08]{e7} Bangert, V; Katz, M.; Shnider, S.; Weinberger,
S.: $E_7$, Wirtinger inequalities, Cayley $4$-form, and homotopy.
{\em Duke Math. J.}, to appear.  See arXiv:math.DG/0608006


\bibitem[Ber76]{Ber} Berstein, I: On the Lusternik--Schnirelmann
category of Grassmannians.  {\em Math. Proc. Camb. Philos. Soc.}
\textbf{79} (1976) 129--134.

\bibitem[BG61]{BG} Berstein, I.; Ganea, T.: The category of a map and
of a cohomology class.  {\em Fund. Math.} \textbf{50} (1961/1962)
265--279.

\bibitem[Bre97]{Br} Bredon, G.: Sheaf Theory. \emph{Graduate Text in
Mathematics}, \textbf{170}, Springer, New York Heidelberg Berlin,
1997.

\bibitem[Bro94]{Bro} Brown, K.: Cohomology of groups. \emph{Graduate
Texts in Mathematics}, \textbf{87} Springer, New York Heidelberg
Berlin, 1994.


\bibitem[CLOT03]{CLOT} Cornea, O.; Lupton, G.; Oprea, J.; Tanr\'e, D.:
Lusternik-Schnirelmann category.  {\em Mathematical Surveys and
Monographs}, \textbf{103}.  American Mathematical Society, Providence,
RI, 2003.

\bibitem[Dr07]{Dr} Dranishnikov, A.: On the Lusternik-Schnirelmann category
of spaces with 2-dimensional fundamental group. arXiv:0709.4018.

\bibitem[DKR08]{DKR2} Dranishnikov, A., Katz, M., Rudyak, Yu.: Small
values of the systolic category of manifolds. Preprint 2008.

\bibitem[DR07]{DR} Dranishnikov, A., Rudyak, Yu.: On the
Berstein-Svarc Theorem in dimension 2.  See arXiv:0712.2087.

\bibitem[Ei47]{E} Eilenberg, S.: Homology of spaces with operators. I.
{\em Trans. Amer. Math. Soc.} \textbf{61} (1947), 378--417; errata,
\textbf{62} (1947), 548.

\bibitem[EG57]{EG} Eilenberg, S; Ganea, T.: On the
Lusternik-Schnirelmann category of abstract groups.  {\em Ann. of
Math.} (2) \textbf{65} (1957), 517--518.

\bibitem[FH92]{FH} Fadell, E.; Husseini, S.: Category weight and
Steenrod operations.  {\em Bol. Soc. Mat. Mexicana\/} (2) \textbf{37}
(1992) no. 1-2, 151--161.

\bibitem[Fe53]{Fe} Fet, A. I.: A connection between the topological
properties and the number of extremals on a manifold (Russian), {\em
Doklady AN SSSR}, \textbf{88} (1953), 415--417.

\bibitem[Fo41]{F} Fox, R.: On the Lusternik--Schnirelmann category,
{\em Ann. of Math.} \textbf{42} (1941) 333--370.



\bibitem[Gr83]{Gr1} Gromov, M.: Filling Riemannian manifolds, {\em
J. Diff. Geom.} \textbf{18} (1983), 1-147.

\bibitem[Gr96]{Gr2} Gromov, M.: Systoles and intersystolic
inequalities, Actes de la Table Ronde de G\'{e}om\'{e}trie
Diff\'{e}rentielle (Luminy, 1992), 291--362, {\em S\'{e}min. Congr.},
\textbf{1}, Soc. Math. France, Paris, 1996.  See
\newline\noindent
www.emis.de/journals/SC/1996/1/ps/smf\_sem-cong\_1\_291-362.ps.gz

\bibitem[Gr99]{Gr3} Gromov, M.: Metric structures for Riemannian and
non-Riemannian spaces, {\em Progr. Math.} \textbf{152},
Birkh\"{a}user, Boston, 1999.


\bibitem[Gr07]{Gr4} Gromov, M.: Metric structures for Riemannian and
non-Riemannian spaces.  Based on the 1981 French original. With
appendices by M. Katz, P. Pansu and S. Semmes. Translated from the
French by Sean Michael Bates. Reprint of the 2001 English
edition. Modern Birkh\"auser Classics.  Birkh\"auser Boston, Inc.,
Boston, MA, 2007




\bibitem[GG92]{GoGo} Gomez-Larra\~naga, J.C.; Gonzalez-Acu\~na, F.:
Lusternik-Schnirelmann category of~$3$-manifolds, {\em Topology\/}
\textbf{31} (1992), 791-800.


\bibitem[H73]{H} Hardie, K.:
On the category of the double mapping cylinder,
\emph{Tohoku Math. J., II. Ser.}\textbf{25} (1973), 355-358.

\bibitem[Ha02]{Ha} Hatcher, A.: Algebraic Topology, Cambridge
University Press, 2002.

\bibitem[Hi04]{Hi} Hillman, J.A. ${\rm PD}\sb 4$-complexes with free
fundamental group.  {\em Hiroshima Math. J.}  \textbf{34} (2004),
no. 3, 295--306.

\bibitem[Ka07]{SGT} Katz, M.: Systolic geometry and topology.  With an
appendix by Jake P. Solomon.  {\em Mathematical Surveys and
Monographs}, \textbf{137}.  American Mathematical Society, Providence,
RI, 2007.




\bibitem[KR06]{KR1} Katz, M.; Rudyak, Y.: Lusternik-Schnirelmann
category and systolic category of low dimensional manifolds.  {\em
Comm. Pure Appl. Math.} \textbf{59} (2006), 1433-1456.  See
arXiv:math.DG/0410456

\bibitem[KR05]{KR2} Katz, M.; Rudyak, Y.: Bounding volume by systoles
of 3-manifolds.  See arXiv:math.DG/0504008

\bibitem[KRS06]{KRS} Katz, M.; Rudyak, Y.; Sabourau, S.:
Systoles of 2-complexes, Reeb graph, and Grushko decomposition.
\emph{Int. Math. Res. Not.} 2006, No. 17, Article ID 54936, 30 p.




\bibitem[OR01]{OR} Oprea, J.; Rudyak, Y.: Detecting elements and
Lusternik-Schnirelmann category of 3-manifolds.
Lusternik-Schnirelmann category and related topics (South Hadley, MA,
2001), 181--191, {\em Contemp. Math.}, \textbf{316}, Amer. Math. Soc.,
Providence, RI, 2002.

\bibitem[Ru97]{R1} Rudyak, Yu.: Category weight: new ideas concerning
Lusternik-Schni\-rel\-mann category.  Homotopy and geometry (Warsaw,
1997), 47--61, Banach Center Publ., 45, Polish Acad. Sci., Warsaw,
1998.

\bibitem[Ru99]{R2} Rudyak, Yu.: On category weight and its
applications.  \emph{Topology} {\bf 38} (1999) no. 1, 37--55.

\bibitem[Sp66]{Sp} Spanier, E.: Algebraic Topology, McGraw-Hill, New
York, 1966.

\bibitem[Sta68]{Stal} Stallings, J.: Groups of dimension 1 are locally
free. \emph{Bull. Amer. Math. Soc.} \textbf{74} (1968), 361--364.

\bibitem[Ste43]{Ste} Steenrod, N.: Homology with local coefficients.
\emph{Ann. of Math.} (2) \textbf{44}, (1943). 610--627.

\bibitem[Str97]{S1} Strom, J.: Category weight and essential category
weight, Thesis, Univ. of Wisconsin 1997.

\bibitem[Str07]{S2} Strom, J.: Lusternik--Schnirelamann Category of
Spaces with Free Fundamental Group, preprint.

\bibitem[Sva66]{Sv} \v Svarc, A.: The genus of a fiber space,
\emph{Amer. Math. Soc. Transl.} Series~2, \textbf{55} (1966), 49--140.

\bibitem[Swan69]{Swan} Swan, R.: Groups of cohomological dimension
one.  {\em J. Algebra\/} \textbf{12} (1969), 585--610.


\bibitem[Wh78]{Wh} Whitehead, G.: Elements of Homotopy
Theory. \emph{Graduate Text in Mathematics}, \textbf{61}, Springer,
New York Heidelberg Berlin, 1978.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb,latexsym, amsmath, amscd, array, graphicx}

\swapnumbers
\numberwithin{equation}{section}

\theoremstyle{plain}

\newtheorem{thm}{Theorem}[section]
\newtheorem{theorem}[thm]{Theorem}

\newtheorem{lemma}[thm]{Lemma}

\newtheorem{prop}[thm]{Proposition}
\newtheorem{cor}[thm]{Corollary}

\newcommand\theoref{Theorem~\ref}
\newcommand\lemref{Lemma~\ref}
\newcommand\propref{Proposition~\ref}
\newcommand\corref{Corollary~\ref}
\newcommand\defref{Definition~\ref}

\newcommand\Zp{\Z_p}

\theoremstyle{definition}

\newtheorem{definition}[thm]{Definition}

\newtheorem{rem}[thm]{Remark}

\newtheorem{ex}[thm]{Example}

\newtheorem{question}[thm]{Question}

\DeclareMathOperator{\sys}{{\rm sys}}

\DeclareMathOperator{\cat}{{\mbox{\rm cat$_{\rm LS}$}}}

\def\Im{\protect\operatorname{Im}}

\def\Hom{\protect\operatorname{Hom}}

\def\pt{\protect\operatorname{pt}}

\def\wgt{\protect\operatorname{wgt}}
\def\cwgt{\protect\operatorname{cwgt}}

\def\cd{\protect\operatorname{cd}}
\def\ber{\mathfrak b}

\newcommand \pa[2]{\frac{\partial #1}{\partial #2}}

\def\gf{\varphi}
\def\ga{\alpha}

\def\scr{\mathcal}
\def\A{{\scr A}}
\def\B{{\scr B}}
\def\O{{\scr O}}

\def\Z{{\mathbb Z}}

\def\R{{\mathbb R}}

\def\zp{\Z[\pi]}

\def\1{\hbox{\rm\rlap {1}\hskip.03in{\rom I}}}
\def\Bbbone{{\rm1\mathchoice{\kern-0.25em}{\kern-0.25em}
{\kern-0.2em}{\kern-0.2em}I}}

\def\pa{\partial}

\def\wt{\widetilde}

\def\m{\medskip}
\def\ov{\overline}

\long\def\forget#1\forgotten{} %

\newcommand\ver[1]{\marginpar{\tiny Changed in Ver \VER}}

\date{\today}
\begin{document}

\title[Small values of LS category for manifolds]{Small values of the
Lusternik-Schnirelmann category for manifolds}

\author[A.~Dranishnikov]{Alexander N. Dranishnikov$^{1}$} %
\thanks{$^{1}$Supported by NSF, grant DMS-0604494}

\author[M.~Katz]{Mikhail G. Katz$^{2}$} %
\thanks{$^{2}$Supported by the Israel Science Foundation (grants 84/03
and 1294/06) and the BSF (grant 2006393)}

\author[Yu.~Rudyak]{Yuli B. Rudyak$^{3}$}%
\thanks{$^{3}$Supported by NSF, grant 0406311}

\address{Alexander N. Dranishnikov, Department of Mathematics, University
of Florida, 358 Little Hall, Gainesville, FL 32611-8105, USA}
\email{dranish@math.ufl.edu}

\address{Mikhail G. Katz, Department of Mathematics, Bar Ilan University,
Ramat Gan 52900 Israel}
\email{katzmik ``at'' math.biu.ac.il}

\address{Yuli B. Rudyak, Department of Mathematics, University
of Florida, 358 Little Hall, Gainesville, FL 32611-8105, USA}
\email{rudyak@math.ufl.edu}

\subjclass[2000]
{Primary 55M30; %LS
Secondary 53C23,  %% Global topological methods (\`a la Gromov)
57N65  %% Algebraic topology of manifolds
}

\keywords{Category weight, cohomological dimension, detecting element,
essential manifolds, free fundamental group, Lusternik-Schnirelmann
category, Massey product, self-linking class, systolic category}

\begin{abstract}
We prove that manifolds of Lusternik-Schnirelmann category~$2$
necessarily have free fundamental group.  We thus settle a 1992
conjecture of Gomez-Larra\~naga and Gonzalez-Acu\~na, by generalizing
their result in dimension~$3$, to all higher dimensions.  We also
obtain some general results on the relations between the fundamental
group of a closed manifold~$M$, the dimension of~$M$, and the
Lusternik-Schnirelmann category of~$M$, and relate the latter to the
systolic category of~$M$.
\end{abstract}

\maketitle
\tableofcontents

\section{Introduction}

We follow the normalization of the Lusternik-Schnirelmann category (LS
category) used in the recent monograph \cite{CLOT} (see
Section~\ref{three} for a definition).  Spaces of LS category~$0$ are
contractible, while a closed manifold of LS category~$1$ is homotopy
equivalent (and hence homeomorphic) to a sphere.

The characterization of closed manifolds of LS category~$2$ was
initiated in 1992 by J.~Gomez-Larra\~naga and F.~Gonzalez-Acu\~na
\cite{GoGo} (see also~\cite{OR}), who proved the following result on
closed manifolds~$M$ of dimension~$3$: the fundamental group of~$M$ is
free and nontrivial if and only if its LS category is~$2$.
Furthermore, they conjectured that the fundamental group of every
closed~$n$-manifold,~$n\ge 3$, of LS category~$2$ is necessarily
free~\cite[Remark, p.~797]{GoGo}.  Our interest in this natural
problem was also stimulated in part by recent work on the comparison
of the LS category and the systolic category~\cite{KR1, KR2, SGT},
which was inspired, in turn, by M.~Gromov's systolic inequalities
\cite{Gr1, Gr2, Gr3, Gr4}.

In the present text we prove this 1992 conjecture.  Recall that all
closed surfaces different from~$S^2$ are of LS category 2.

\begin{theorem}
\label{11}
A closed connected manifold of LS category~$2$ either is a surface, or
has free fundamental group.
\end{theorem}

\begin{cor}
\label{12b}
Every manifold~$M^n, n\geq 3$, with non-free fundamental group
satisfies~$\cat(M)\geq 3$.
\end{cor}

We found that there is no restriction on the fundamental group
for closed manifolds of LS category 3. In particular we proved
the following.
\begin{theorem}
\label{t:realization}
Given a finitely presented group~$\pi$ and non-negative integers~$k,
l$, there exists a closed manifold~$M$ such that~$\pi_1(M)=\pi$,
while~$\cat M=3+k$ and~$\dim M=5+2k+l$.  Furthermore, if~$\pi$ is not
free, then~$M$ can be chosen~$4$-dimensional with~$\cat M=3$.
\end{theorem}

Thus, there is no restriction on the fundamental group of manifolds of
LS category 3 and higher.

The above results lead to the following questions:

\begin{question}
\label{4-problem}
If a 4-dimensional CW-complex ~$X$ has free fundamental group, then we
have the bound~$\cat X \leq 3$.  Is the stronger bound~$\cat X \leq 2$
necessarily satisfied?
\end{question}

We prove the inequality $\cat M \le n-2$ for connected~$n$-manifolds
with free fundamental group and~$n>4$, see \propref{p:n>4}.
In~\cite{S2}, J.~Strom proved a stronger inequality~$\cat X\le
\frac{2}{3}\dim X$ for an arbitrary CW-space $X$.  Later, it was
proved in~\cite{Dr} that if the fundamental group is free, then the
bound
\begin{equation}
\cat X\le \frac{1}{2}\dim X +1,
\end{equation}
is satisfied by every CW-complex~$X$.

The above Question~\ref{4-problem} has an affirmative answer when~$M$
is a closed orientable manifold, in view of a theorem due to
J.~A.~Hillman~\cite{Hi} which states that a closed~$4$-dimensional
manifold with free fundamental group has a CW-decomposition in which
the three-skeleton has the homotopy type of a wedge of spheres.

\begin{question} 
Is it true that~$\cat (M \setminus \{\pt\})=1$ for any closed
manifold~$M$ with~$\cat M=2$?  This is proved in~\cite{GoGo} for the
case~$\dim M=3$. A direct proof would imply the main theorem
trivially.
\end{question}

\begin{question}
Given integers~$m$ and~$n$, describe the fundamental groups of closed
manifolds~$M$ with~$\dim M=n$ and~$\cat M=m$.
\end{question}

Note that in the case~$m=n$, the fundamental group of~$M$ is of
cohomological dimension~$\ge n$, see e.g. the Berstein--\v Svarc
\theoref{t:berstein}.
Thus, we can ask when the converse holds.

\begin{question}
Given a finitely presented group~$\pi$ and an integer $n\geq 4$ such
that~$H^n(\pi)\not = 0$, when can one find a closed manifold~$M$
satisfying~$\pi_1(M)=\pi$ and $\dim M=\cat M=n?$ Note that
\propref{p:lense} shows that such a manifold~$M$ does not always
exist.
\end{question}

A related numerical invariant called the {\em systolic category\/} can
be thought of as a Riemannian analogue of the LS category~\cite{SGT}.
In \cite{DKR2} we apply \corref{12b} to prove that the systolic
category of a~$4$-manifold is a lower bound for its LS category.

\begin{theorem}
\label{12}
Every closed orientable~$4$-manifold~$M$ satisfies the
inequality~${\rm cat}_{\sys}(M) \leq \cat(M)$.
\end{theorem}

In particular, this inequality implies that if a 4-manifold $M$ has a
free fundamental group then ${\rm cat}_{\sys}(M) = \cat(M)$.  In a
related development in systolic topology, an intriguing model for
$BS^3$ built out of $BS^1$ was used in \cite{e7} to prove that the
symmetric metric of the quaternionic projective space, contrary to
expectation, is {\em not\/} its systolically optimal metric.

\m The proof of the main theorem proceeds roughly as follows. If the
group~$\pi:=\pi_1(M)$ is not free, then by a result of J.~Stallings
and R.~Swan, the group~$\pi$ is of cohomological dimension at
least~$2$.  We then show that~$\pi$ carries a suitable
nontrivial~$2$-dimensional cohomology class~$u$ with twisted
coefficients, and of category weight~$2$.  Viewing~$M$ as a subspace
of~$K(\pi,1)$ that contains the~$2$-skeleton~$K(\pi,1)^{(2)}$, and
keeping in mind the fact that the~$2$-skeleton carries the fundamental
group, we conclude that the restriction (pullback) of~$u$ to~$M$ is
non-zero and also has category weight~$2$.  By Poincar\'e duality with
twisted coefficients, one can find a complementary~$(n-2)$-dimensional
cohomology class.  By a category weight version of the cuplength
argument, we therefore obtain a lower bound of~$3$ for~$\cat M$.

In Section~\ref{local}, we review the material on local coefficient
systems, a twisted version of Poincar\'e duality, and 2-dimensional
cohomology of non-free groups.  In Section~\ref{three}, we review the
notion of category weight.  In Section~\ref{four}, we prove our main
result, Theorem~\ref{11}.  In Section~\ref{higher} we prove
\theoref{t:realization}.

\section{Cohomology with local coefficients}
\label{local}

A {\it local coefficient system}~$\A$ on a path
connected~CW-space~$X$ is a functor from the fundamental
groupoid~$\Gamma(X)$ of~$X$, to the category of abelian groups.  See
\cite{Ha}, \cite{Wh} for the definition and properties of local
coefficient systems.

In other words, an abelian group~$\A_x$ is assigned to each
point~$x\in X$, and for each path~$\ga$ joining~$x$ to~$y$, an
isomorphism~$\ga^*:\A_y \to \A_x$ is given.  Furthermore, paths that
are homotopic are required to yield the same isomorphism.

Let~$\pi=\pi_1(X)$, and let~$\zp$ be the group ring of~$\pi$.  Note
that all the groups~$A_x$ are isomorphic to a fixed group~$A$.  We
will refer to~$A$ as a {\it stalk\/} of~$\A$.

Given a map~$f: Y \to X$ and a local coefficient system~$\A$ on~$X$,
we define a local coefficient system on~$Y$, denoted~$f^*\A$, as
follows.  The map~$f$ yields a functor~$\Gamma(f): \Gamma(Y) \to
\Gamma(X)$, and we define~$f^*\A$ to be the functor~$\A\circ
\Gamma(f)$. Given a pair of coefficient systems~$\A$ and~$\B$, the
tensor product~$\A\otimes \B$ is defined by setting~$(\A\otimes
\B)_x=\A_x\otimes\B_x$.

\m \begin{ex}\label{ex:bundle} A useful example of a local coefficient
system is given by the following construction.  Given a fiber bundle
$p: E \to X$ over~$X$, set~$F_x=p^{-1}(x)$.  Then the family
$\{H_k(F_x)\}$ can be regarded a local coefficient system, see
\cite[Example 3, Ch. VI, \S 1]{Wh}.  An important special case is that
of an~$n$-manifold~$M$ and spherical tangent bundle~$p: E \to M$ with
fiber~$S^{n-1}$, yielding a local coefficient system~$\O$ with
$\O_x=H_{n-1}(S^{n-1}_x)\cong \Z$.  This local system is called the
{\it orientation sheaf\/} of~$M$.
\end{ex}

\begin{rem}
\label{22}
There is a bijection between local coefficients on~$X$ and
$\zp$-modules \cite[Ch. 1, Exercises F]{Sp}. If~$\A$ is a local
coefficient system with stalk~$A$, then the natural action of the
fundamental group on~$A$ turns~$A$ into a~$\Z[\pi]$-module.
Conversely, given a~$\Z[\pi]$-module~$A$, one can construct a local
coefficient system~$\scr L(A)$ such that induced~$\Z[\pi]$-module
structure on~$A$ coincides with the given one, cf.~\cite{Ha}.
\end{rem}

We recall the definition of the (co)homology groups with local
coefficients via modules \cite{Ha}:
\begin{equation}\label{cohomology}
H^k(X;\A)\cong H^k(\Hom_{\zp}(C_*(\wt X), A), \delta)
\end{equation}
and
\begin{equation}\label{homology}
H_k(X;\A)\cong H_k(A\otimes _{\zp}C_*(\wt X), 1\otimes \pa).
\end{equation}
Here~$(C_*(\wt X), \pa)$ is the chain complex of the universal
cover~$\wt X$ of~$X$,~$A$ is the stalk of the local coefficient
system~$\A$, and~$\delta$ is the coboundary operator.  Note that in
the tensor product we used the right~$\zp$ module structure on~$A$
defined via the standard rule~$ag=g^{-1}a$, for~$a\in A, g\in \pi$.

Recall that for~CW-complexes~$X$, there is a natural bijection
between equivalence classes of local coefficient systems and locally
constant sheaves on~$X$.  One can therefore define (co)homology with
local coefficients as the corresponding sheaf cohomology \cite{Br}.
In particular, we refer to \cite{Br} for the definition of the cup
product
\[
\cup: H^i(X;\A) \otimes H^j(X;\B) \to H^{i+j}(X; \A\otimes \B)
\]
and the cap product
\[
\cap: H_i(X;\A) \otimes H^j(X;\B) \to H_{i-j}(X; \A\otimes \B).
\]
A nice exposition of the cup  and the cap products in a slightly
different setting can be found in \cite{Bro}.  In particular, we have
the cap product
\[
H_k(X;\A)\otimes H^k(X;\B)\to H_0(X;\A\otimes \B)\cong
A\otimes_{\zp}B.
\]

\begin{prop}
\label{l:evaluation}
Given an integer~$k\ge 0$, there exists a local coefficient
system~$\B$ and a class~$v\in H^k(X;\B)$ such that, for every local
coefficient system~$\A$ and nonzero class~$a\in H_k(X;\A)$, we
have~$a\cap v\ne 0$.
\end{prop}

\begin{proof}
Throughout the proof $\otimes$ denotes $\otimes_{\zp}$. We convert the
stalk of~$\A$ into a right~$\zp$-module~$A$ as above.  Below we use the
isomorphisms~\eqref{cohomology} and \eqref{homology}.  Consider the
chain~$\zp$-complex
\[
\CD \dots @>>> C_{k+1}(\wt X) @>\pa_{k+1}>>C_k(\wt X) @>\pa_k>>
C_{k-1}(\wt X) @>>> \dots .
\endCD
\]
For the given~$k$, we set~$B: =C_k(\wt X)/\Im \pa_{k+1}$.  Let~$\B$ be
the corresponding local system on~$X$.  Thus, we obtain the exact
sequence of~$\zp$-modules
\[
\CD C_{k+1}(\wt X) @>\pa_{k+1}>>C_k(\wt X)@>f>>B\to 0.  \endCD
\]
Note that the epimorphism~$f$ can be regarded as a~$k$-cocycle with
values in~$\B$, since~$\delta f (x)=f\pa_{k+1}(x)=0$.  Let~$v:=[f]\in
H^k(X;\B)$ be the cohomology class of~$f$.  Now we prove that
$$
a\cap [f] \not=0.
$$

Since the tensor product is right exact, we obtain the diagram
\[
\CD A\otimes C_{k+1}(\wt X)
@>1\otimes\pa_{k+1}>>A\otimes C_k(\wt X)@>1\otimes
f>>A\otimes B @>>> 0\\ @. @. g @ VVV @.\\
@. @. A\otimes C_{k-1}(\wt X) \endCD
\]
where the row is exact.  The composition
\[
\CD A\otimes C_k(\wt X)@>1\otimes f>>A\otimes B @>g>>
A\otimes C_{k-1}(\wt X)
\endCD
\]
coincides with~$1\otimes \pa_k$. We represent the class~$a$ by a cycle
\[
z\in A\otimes  C_k(\wt X).
\]
Since~$z\notin \Im(1\otimes \pa_{k+1})$, we conclude that
$$
(1\otimes f)(z)\ne 0\in A\otimes B=H_0(X;\A\otimes \B).
$$
Thus, for the cohomology class~$v$ of~$f$ we have~$a\cap v \ne 0$.
\end{proof}

Every closed connected~$n$-manifold~$M$ satisfies~$H_n(M; \O)\cong
\Z$.  A generator (one of two) of this group is called the {\it
fundamental class} of~$M$ and is denoted by~$[M]$.

\m One has the following generalization of the Poincar\'e duality
isomorphism.

\begin{thm}[{\cite[Corollary 10.2]{Br}}]
\label{th:PD}
The homomorphism
\begin{equation}\label{eq:PD}
\Delta: H^i(M;\A)\to H_{n-i}(M;\O\otimes \A)
\end{equation}
defined by setting~$\Delta(a)=[M]\cap a$, is an isomorphism.
\end{thm}

In fact, in \cite{Br} there is the sheaf~$\O^{-1}$ at the right, but
for manifolds we have~$\O=\O^{-1}$.

Given a group~$\pi$ and a~$\Z[\pi]$-module~$A$, we denote by
$H^*(\pi;A)$ the cohomology of the group~$\pi$ with coefficients
in~$A$, see e.g.~\cite{Bro}.  Recall that~$H^i(\pi;A)=
H^i(K(\pi,1);\scr L(A))$, see Remark~\ref{22}.

\m Let~$\cd(\pi)$ denote the cohomological dimension of~$\pi$
over~$\Z$, i.e. the largest~$m$ such that there exists
an~$\Z[\pi]$-module~$A$ with~$H^m(\pi; A)\ne 0$.

\begin{thm}[\cite{Stal, Swan}]\label{th:free}
If~$\cd\pi\le 1$ then~$\pi$ is a free group.
\end{thm}

We will need the following known fact from the cohomology theory of
groups.

\begin{lemma}\label{l:cd}
If~$\pi$ be a group with~$\cd\pi=q\geq 2$.  Then~$H^2(\pi;A)\ne
0$ for some~$\zp$-module~$A$.
\end{lemma}

\begin{proof} 
We use the fact that cohomology of the group $\pi$ with coefficients
in an injective $\zp$-module are trivial and the fact that every
$\zp$-module $A'$ can be imbedded into an injective $\zp$module
$J$,~\cite{Bro}.  Let~$0\to A'\to J\to A''\to 0$ be an exact sequence
of~$\zp$-modules with~$J$ injective. Then by the coefficients long
exact sequence~$H^k(\pi;A')=H^{k-1}(\pi;A'')$ for $k>1$.
Since~$H^q(\pi;B)\ne 0$ for some~$B$, the proof can be completed by an
obvious induction.
\end{proof}

\section{Category weight and lower bounds for~$\cat$}
\label{three}

In this section, we review the notion of category weight and its
relation to the Lusternik-Schnirelmann category.

\begin{definition}[\cite{BG,Fe,F}]
\label{def:cat-map}
Let~$f: X \to Y$ be a map of (locally contractible)~CW-spaces. The
\emph{Lusternik--Schnirelmann category of~$f$}, denoted~$\cat(f)$, is
defined to be the minimal integer~$k$ such that there exists an open
covering~$\{U_0, \ldots, U_k\}$ of~$X$ with the property that each of
the restrictions~$f|A_i\colon A_i \to Y$,~$i=0,1, \ldots, k$ is
null-homotopic.

The \emph{Lusternik--Schnirelmann category~$\cat X$ of a space~$X$} is
defined as the category~$\cat (1_X)$ of the identity map.
\end{definition}

\begin{definition}\label{def:swgt}
The \emph{category weight}~$\wgt(u)$ of a non-zero cohomology class~$u \in
H^*(X; \A)$ is defined as follows:
\begin{equation*}
\label{31}
\wgt(u)\ge k \Longleftrightarrow \{\gf^*(u)=0 {\rm\ for\ every\ }
\gf\colon F \to X {\rm\ with\ } \cat(\gf) < k\}.
\end{equation*}
\end{definition}

\begin{rem}\label{rem:credits}\rm
E.~Fadell and S.~Husseini (see \cite{FH}) originally proposed the
notion of category weight.  In fact, they considered an invariant
similar to the~$\wgt$ of \eqref{31} (denoted in \cite{FH} by~$\cwgt$),
but where the defining maps~$\gf\colon F \to X$ were required to be
inclusions rather than general maps.  As a consequence,~$\cwgt$ is not
a homotopy invariant, and thus a delicate quantity in homotopy
calculations.  Yu.~Rudyak \cite{R1, R2} and J.~Strom \cite{S1}
proposed a homotopy invariant version of category weight as defined in
\defref{def:swgt}.
\end{rem}

\begin{prop}[\cite{R1,S1}]
\label{prop:swgtprops}
Category weight has the following properties.
\begin{enumerate}
\item~$1\le \wgt(u) \leq \cat(X)$, for all~$u \in \widetilde
H^*(X;\A), u\ne 0$.
\vskip3pt
\item For every~$f\colon Y \to X$ and~$u\in H^*(X;\A)$ with
$f^*(u)\not = 0$ we have
$\cat(f) \geq \wgt(u)$ and~$\wgt(f^*(u)) \geq \wgt(u)$.
\vskip3pt
\item For~$u\in H^*(X;\A)$ and~$v\in H^*(X;\B)$ we have
\[
\wgt(u\cup v) \geq \wgt(u) + \wgt(v).
\]  \vskip3pt
\item For every~$u \in H^s(K(\pi,1);\A)$,~$u\ne 0$, we have
$\wgt(u)\geq s$.  \vskip3pt
\end{enumerate}
\end{prop}

\begin{proof}
See \cite[\S 2.7 and Proposition 8.22]{CLOT}, the proofs in
loc. cit. can be easily adapted to local coefficient systems.
\end{proof}

\section{Manifolds of LS category~$2$}
\label{four}

In this section we prove that the fundamental group of a closed
connected manifold of LS category~$2$ is free.

\begin{thm}
\label{th:main}
Let~$M$ be a closed connected manifold of dimension at least~$3$.  If
the group~$\pi:=\pi_1(M)$ is not free, then~$\cat M \geq 3$.
\end{thm}

\begin{proof}
By \theoref{th:free} and \lemref{l:cd}, there a local coefficient
system~$\A$ on~$K(\pi,1)$ such that~$H^2(K(\pi,1);\A)\ne 0$.  Choose a
non-zero element~$u\in H^2(K(\pi,1);\A)$. Let~$f: M \to K(\pi,1)$ be
the map that induces an isomorphism of fundamental groups, and let~$i:
K\to M$ be the inclusion of the~$2$-skeleton. (If~$M$ is not
triangulable, we take~$i$ to be any map of a~$2$-polyhedron that
induces an isomorphism of fundamental groups.)  Then
\[
(fi)^*: H^2(K(\pi,1);\A) \to H^2(K;(fi)^*\A)
\]
is a monomorphism.  In particular, we have~$f^*u\not=0$
in~$H^2(M;(f)^*\A)$.  Now consider the class
\[
a=[M]\cap f^*u\in H_{n-2}(M;\O^{-1}\otimes f^*\A),
\]
where $n=\dim M$.  Then~$a\ne 0$ by \theoref{th:PD}.  Hence, by
Proposition~\ref{l:evaluation}, there exists a class~$v\in
H^{n-2}(M;\B)$ such that~$a\cap v \ne 0$. We claim that~$f^*u\cup v
\ne 0$. Indeed, one has
\[
[M]\cap (f^*u\cup v)=([M]\cap f^*u)\cap v =a\cap v\ne 0.
\]
Now,~$\wgt f^*u\ge 2$ by \propref{prop:swgtprops}, items (2)
\and~(4). Furthermore,~$\wgt(v)\geq 1$ by \propref{prop:swgtprops},
item~(1).  We therefore obtain the lower bound~$\wgt(f^*u\cup v)\ge 3$
by \propref{prop:swgtprops}, item (3).  Since~$f^*u\cup v\ne 0$, we
conclude that~$\cat M \ge 3$ by \propref{prop:swgtprops}, item (1).
\end{proof}

\begin{cor}
If~$M^n, n\geq 3$ is a closed manifold with~$\cat M\le 2$, then
$\pi_1(M)$ is a free group.
\end{cor}

\begin{rem}
An alternative approach to Theorem~\ref{th:main} would be using the
Berstein-\v{S}varc class $\ber\in H^1(\pi;I(\pi))$ where $I(\pi)$ is
the augmentation ideal of $\pi$. If~$ \cd(\pi)\ge 2$ then $\ber^{2}
\ne 0$ by \cite{DR} (see also \theoref{t:berstein}).  In
particular,~$H^2(\pi;I(\pi)\otimes I(\pi))\ne 0$, and we obtain an
alternative proof of \lemref{l:cd}.
\end{rem}

The following Proposition is a special case of ~\cite[Corollary
4.2]{Dr}.  Here we give a relatively simple geometric proof.

\begin{prop}\label{p:n>4}
Let $M$ be a closed connected $n$-dimensional PL manifold, $n>4$, with
free fundamental group.  Then $\cat M \le n-2$.
\end{prop}

\begin{proof}
If $X$ is a 2-dimensional (connected) CW-complex with free fundamental
group then $\cat X\le 1$, see e.g.~\cite[Theorem 12.1]{KRS}.  Hence,
if $Y$ is a $k$-dimensional complex with free fundamental group then
$\cat Y\le k-1$ for $k>2$.  Now, let $K$ be a triangulation of $M$,
and let $L$ be its dual triangulation.  Then $M\setminus L^{(l)}$ is
homotopy equivalent to $K^{(k)}$ whenever $k+l+1=n$.  Hence,
\[
\cat M\le \cat K^{(k)}+\cat L^{(l)}+1.
\]
Since $\pi_1(K)$ and $\pi_1(L)$ are free, we conclude that $\cat
K^{(k)}\le k-1$ and $\cat L^l\le l-1$ for $k,l>1$. Thus $\cat M\le k-1
+l-1 +1=n-2$.
\end{proof}

\section{Manifolds of higher LS category}
\label{higher}

Gromov~\cite[4.40]{Gr3} called a polyhedron $X$ {\em $n$-essential} if
there is no map $f:X\to K(\pi,1)^{(n-1)}$ to the $(n-1)$-dimensional
skeleton of an Eilenberg-MacLane complex that induces an isomorphism
of the fundamental groups. We extend his definition as follows.

\begin{definition}
\label{def:k-essent}
A~CW-space~$X$ is called {\em strictly $k$-essential},~$k>1$ if for
every CW-complex structure on~$X$ there is no map between the skeleta
$f:X^{(k)}\to K(\pi,1)^{(k-1)}$ that induces an isomorphism of the
fundamental groups.
\end{definition}

Clearly, a strictly $n$-essential space is Gromov $n$-essential, while
the converse is false. Furthermore, an $n$-dimensional polyhedron is
strictly $n$-essential if it is Gromov $n$-essential.

\begin{thm}\label{th:k-essent}
Let $M$ be a closed strictly ~$k$-essential manifold.  If its
dimension satisfies~$\dim M \geq k+1$, then its LS category also
satisfies $\cat M\ge k+1$.
\end{thm}

\begin{proof}
We first consider the case~$k=2$. If~$\cat M\le 2$, then, by
\theoref{th:main},~$\pi_1(M)$ is free.  Hence there is a map~$f:M\to
\vee S^1$ that induces an isomorphism of the fundamental groups,
and~$M$ is not strictly $2$-essential.

Now assume~$k\ge 3$.  Let~$K=K(\pi_1(M),1)$.  Consider a map
\[
f: M^{(k-1)} \to K^{(k-1)}
\]
such that the restriction~$f|_{M^{(2)}}$ is the identity homeomorphism
of the 2-skeleta~$M^{(2)}$ and~$K^{(2)}$. We consider the problem of
extension of~$f$ to~$M$.

We claim that the first obstruction ~$o(f)\in H^k(M;E)$ (taken with
coefficients in a local system~$E$ with the stalk
$\pi_{k-1}(K^{(k-1)})$) to the extension is not equal to zero.

Indeed, if~$o(f)=0$, then there exists a map~$\ov f:M^{(k)}\to
K^{(k-1)}$ which coincides with~$f$ on the~$(k-2)$-skeleton. The
map
\[
\ov f_*: \pi_1(M^{(k)})\to \pi_1(K^{(k-1)})
\]
can be viewed as an endomorphism of~$\pi_1(M)$ that is identical on
generators, and therefore~$\ov f_*$ is an isomorphism.  Hence~$M$ is
not strictly~$k$-essential.

Consider the commutative diagram
\[
\CD
M^{(k-1)} @>f>> K^{(k-1)} @>{\rm id}>> K^{(k-1)}\\
@ViVV @VjVV @.\\
M @>\tilde f>> K
\endCD
\]
where~$i$ and~$j$ are the inclusions of the skeleta.  Let~$\alpha$ be
the first obstruction to the extension of id to a map~$K \to
K^{(k-1)}$.  By commutativity of the above diagram, we
have~$o(f)=\tilde f^*(\alpha)$.  Now, asserting as in the proof of
\theoref{th:main}, we get that~$\tilde f^*(\alpha)\cup v\ne 0$ for
some~$v$ with~$\dim v=\dim M-k$.  Since~$\dim M>k$, we conclude
that~$\dim v\ge 1$ and thus~$\cat M\ge k+1$.
\end{proof}

\begin{rem}
If a closed manifold~$M^n$ is~$n$-essential then~$\cat M=n$, see
e.g.~\cite{KR1} and~\cite[Theorem 12.5.2]{SGT}.
\end{rem}

The following theorem for $n\geq 3$ was proven in~\cite[Theorem
A]{Ber} and~\cite[Theorem 20]{Sv}, see also~\cite[Proposition
2.51]{CLOT}. The case $n=2$ was proved in \cite{DR}.

\begin{theorem}
\label{t:berstein}
If $\dim X=\cat X=n$, then $u^{n}_X\ne 0$ where
$u_X=j^*(\ber)\in H^1(X;I(\pi))$, $j:X\to K(\pi,1)$ induces an isomorphism
of the fundamental groups, and $\ber\in H^1(\pi, I(\pi))$ is the
Berstein-\v{S}varc class.
 {\rm (}For the case
~$n=\infty$ this means that~$u^{k}\ne 0$ for all~$k$.{\rm )}
\end{theorem}

\begin{prop}\label{p:nonfree}
For every non-free finitely presented group~$\pi$, there exists a
closed~$4$-dimensional manifold~$M$ with fundamental group~$\pi$
and~$\cat M=3$.
\end{prop}

\begin{proof}
Let $K$ be a~$2$-skeleton of~$K(\pi,1)$. Take an embedding of $K$ in~$\R^5$
and let~$M=\partial N$ be the boundary of the regular neighborhood
$N$ of this skeleton. Then there is a retraction $N \to K$, and, clearly,
the map $f: M \subset N \to K$ induces an isomorphism of fundamental groups.
Now, let $u_M\in H^1(M; I(\pi))$ be the class described in
the Theorem~\ref{t:berstein}.
Then $u_M=f^*u_K$, and hence $u_M^{4}=0$.
 Therefore~$\cat M<4$ by \theoref{t:berstein}, and thus $\cat M= 3$.
\end{proof}

Let~$M_f$ be the mapping cylinder of~$f:X\to Y$.  We use the notation
$\pi_*(f)=\pi_*(M_f,X)$.  Then~$\pi_i(f)=0$ for~$i\le n$ amounts to
saying that it induces isomorphisms~$f_*:\pi_i(X_1)\to \pi_i(Y_1)$ for
$i\le n$ and an epimorphism in dimension~$n+1$. Similar notation
$H_*(f)=H_*(M_f,X)$ we use for homology.

\begin{lemma}
\label{l:join} Let~$f_j:X_j\to Y_j$ be a family of maps of~CW-spaces such
that~$H_i(f_j)=0$ for~$i\le n_j$. Then~$ H_i(f_1\wedge
\cdots \wedge f_s)=0$ for~$i\le \min\{n_j\}$.
\end{lemma}

\begin{proof}
Note that
\[
M(f_1\wedge \cdots \wedge f_s)\cong Y_1\wedge \cdots \wedge Y_s\cong
M(f_1)\wedge \cdots \wedge M(f_s).
\]
Now, by using the K\"unneth formula and considering the homology exact
sequence of the pair~$(M(f_1)\wedge\cdots \wedge M(f_s), X_1\wedge
\cdots \wedge X_s)$, we obtain the result.
\end{proof}

\begin{prop}\label{p:join}
Let~$f_j:X_j\to Y_j$,~$3\le j\le s$ be a family of maps of~CW-spaces such
that~$\pi_i(f_j)=0$ for~$i\le n_j$.  Then the joins
satisfy
\[
\pi_k(f_1\ast f_2\ast\dots\ast f_s)=0
\]
for~$k\le\min\{n_j\}+s-1$.
\end{prop}

\begin{proof}
By the version of the Relative Hurewicz Theorem for non-simply
connected~$X_j$ \cite[Theorem 4.37]{Ha}, we obtain~$H_i(f_j)=0$ for
$i\le n_j$.  By \lemref{l:join} we obtain
that~$H_k(f_1\wedge\cdots\wedge f_s)=0$ for~$k\le\min\{n_j\}$.  Since
the join~$A_1\ast\dots\ast A_s$ is homotopy equivalent to the iterated
suspension~$\Sigma^{s-1}(A_1\wedge\dots\wedge A_s)$ over the smash
product, we conclude that~$H_k(f_1\ast\dots\ast f_s)=0$
for~$k\le\min\{n_j\}+s-1$.  Since~$X_1\ast\dots\ast X_s$ is simply
connected for~$s\ge 3$, by the standard Relative Hurewicz Theorem we
obtain that~$\pi_k(f_1\ast\dots \ast f_s)=0$ for $k\le
\min\{n_j\}+s-1$.
\end{proof}

Given two maps~$f:Y_1\to X$ and~$g:Y_2\to X$, we set
\[
Z=\{(y_1,y_2,t)\in Y_1\ast Y_2\mid f(y_1)=g(y_2)\}
\]
and define the {\em fiberwise join}, or {\em join over~$X$} of~$f$
and~$g$ as the map
\[
f{\ast_X}g:Z\to X,\quad (f{\ast_X}g)(y_1,y_2, t)=f(y_1)
\]
Let~$p_0^X:PX\to X$ be the Serre path fibration. This means that
$PX$ is the space of paths on~$X$ that start at the base point of the
pointed space~$X$, and~$p_0(\alpha)=\alpha(1)$.  We denote by
$p_n^X;G_n(X)\to X$ the~$n$-fold fiberwise join of~$p_0$.

The proof of the following theorem can be found in \cite{CLOT}.

\begin{thm}[Ganea, \v Svarc]
\label{t:ganea}
For a~CW-space~$X$,~$\cat(X)\le n$ if and only if there exists a
section of~$p_n:G_n(X)\to X$.
\end{thm}

\begin{prop}\label{p:sum}
The connected sum~$S^k\times S^l\#\cdots \#S^k\times S^l$ is a space
of LS-category~$2$.
\end{prop}

\begin{proof}
This can be deduced from a general result of K. Hardy~\cite{H} because
the connected sum of two manifolds can be regarded as the double
mapping cylinder. Alternatively, one can note that, after removing a
point, the manifold on hand is homotopy equivalent to the wedge of
spheres.
\end{proof}

\begin{thm}
\label{t:high}
For every finitely presented group~$\pi$ and~$n\ge 5$, there is a
closed~$n$-manifold~$M$ of LS-category~$3$ with~$\pi_1(M)=\pi$.
\end{thm}

\begin{proof}
If the group~$\pi$ is the free group of rank~$s$, we let~$M'$ be the
$k$-fold connected sum~$S^1\times S^{2} \# \cdots \#S^1\times S^{2}$.
Then~$M'$ is a closed~$3$-manifold of LS category~$2$ with
$\pi_1(M')=F_s$.  Then the product manifold~$M=M'\times S^{n-3}$ has
cuplength~$3$ and is therefore the desired manifold.

Now assume that the group~$\pi$ is not free.  We fix a presentation
of~$\pi$ with~$s$ generators and~$r$ relators.  Let~$M'$ be
the~$k$-fold connected sum~$S^1\times S^{n-1} \# \cdots \#S^1\times
S^{n-1}$.  Then~$M'$ is a closed~$n$-manifold of the category~$2$ with
$\pi_1(M')=F_s$.  For every relator~$w$ we fix a nicely imbedded
circle~$S^1_w\subset M'$ such that~$S_w^{-1}\cap S_v^{-1}=\emptyset$
for~$w\ne v$.  Then we perform the surgery on these circles to obtain
a manifold~$M$. Clearly,~$\pi_1(M)=\pi$. We show that~$\cat(M)\le 3$,
and so~$\cat M =3$ by \theoref{th:main}.

As usual, the surgery process yields an~$(n+1)$-manifold~$X$ with
$\partial X=M\sqcup M'$.  Here~$X$ is the space obtained from
$M'\times I$ by attaching handles~$D^2\times D^{n-1}$ of index 2 to
$M'\times 1$ along the above circles. We note that~$\cat(X)\le 3$.

On the other hand, by duality,~$X$ can be obtained from~$M \times I$
by attaching handles of index~$n-1$ to the boundary component of~$M
\times I$.  In particular, the inclusion~$f: M \to X$ induces an
isomorphism of the homotopy groups of dimension~$\le n-3$ and an
epimorphism in dimension~$n-2$.  Hence the map
\[
\Omega f:\Omega M\to\Omega X
\]
induces isomorphisms in dimensions~$\le n-4$ and an epimorphism in
dimension~$n-3$.  Thus,~$\pi_i(\Omega f)=0$ for~$i\le n-3$.

In order to prove the bound~$\cat M\le 3$, it suffices to show that
the Ganea-\v Svarc fibration~$p_3:G_3(M)\to M$ has a section.
Consider the commutative diagram
\[
\CD
G_3M @>q> >Z @>f'>> G_3(X)\\
@Vp_M^3VV @Vp'VV @ VVp_3^XV\\
M @= M @>f>> X\\
\endCD
\]
where the right-hand square is the pull-back diagram
and~$f'q=G_3(f)$. Note that~$q$ is uniquely determined.
Since~$\cat(X)\le 3$, by \theoref{t:ganea} there is a section~$s:X\to
G_3(X)$.  It defines a section~$s':M\to Z$ of~$p'$. It suffices to
show that the map~$s':M\to Z$ admits a homotopy lifting~$h: M \to
G_3M$ with respect to~$q$, i.e. the map~$h$ with ~$qh\cong
s'$. Indeed, we have
\[
p_M^3h=p'qh\cong p's'=1_M
\]
and so~$h$ is a homotopy section of~$p_3^M$. Since the latter is a
Serre fibration, the homotopy lifting property yields an actual
section.

Let~$F_1$ and~$F_2$ be the fibers of fibrations~$p_3^M$ and~$p'$,
respectively. Consider the commutative diagram generated by the homotopy exact
sequences
of the Serre fibrations~$p_3^M$ and~$p'$:
\[
\CD
\pi_i(F_1) @>>> \pi_i(G_3(M)) @>(p_3^M)_*>> \pi_i(M) @>>>
\pi_{i-1}(F_1) @>>>\cdots\\
 @VV\phi_*V @ VVq_*V @VV=V @VV\phi_*V @.\\
\pi_i(F_2) @>>> \pi_i(Z) @>(p')_*>> \pi_i(M) @>>>\pi_{i-1}(F_2)@>>>
\cdots .\\
\endCD
\]
Note that we have
\[
\phi=\Omega (f)\ast \Omega(f)\ast\Omega(f)\ast \Omega(f).
\]

By \propref{p:join} and since~$\pi_i(\Omega f)=0$ for~$i\le n-3$, we
conclude that~$\pi_i(\phi)=0$ for~$i\le n-3+3=n$.  Hence~$\phi$
induces an isomorphism of the homotopy groups of dimensions~$\le n-1$
and an epimorphism in dimension~$n$.  By the Five Lemma we obtain
that~$q_*$ is an isomorphism in dimensions~$\le n-1$ and an
epimorphism in dimension~$n$. Hence the homotopy fiber of~$q$
is~$(n-1)$-connected. Since~$\dim M=n$, the map~$s'$ admits a homotopy
lifting~$h:M\to G_3(M)$.
\end{proof}

\begin{cor}
\label{c:real}
Given a finitely presented group~$\pi$ and non-negative integer
numbers~$k, l$ there exists a closed manifold~$M$ such
that~$\pi_1(M)=\pi$, while~$\cat M=3+k$ and~$\dim M=5+2k+l$.
\end{cor}

\begin{proof}
By \theoref{t:high}, there exists a manifold~$N$ such that
$\pi_1(M)=\pi$,~$\cat M=3$ and~$\dim M=5+l$. Moreover, this manifold
$N$ possesses a detecting element, i.e. a cohomology class whose
category weight is equal to~$\cat N=3$. For~$\pi$ free this follows
since the cuplength of~$N$ is equal to 3, for other groups we have the
detecting element~$f^*u\cup v$ constructed in the proof of
\theoref{th:main}. If a space~$X$ possesses a detecting element then,
for every~$k>0$, we have~$\cat (X\times S^k) =\cat X+1$ and~$X \times
S$ possesses a detecting element, \cite{R2}. Now, the manifold
$M:=N\times (S^2)^k$ is the desired manifold.
\end{proof}

Generally, we have a question about relations between the category,
the dimension, and the fundamental group of a closed manifold. The
following proposition shows that the situation quite intricate.

\begin{prop}
\label{p:lense}
Let $p$ be an odd prime. Then there exists a closed $(2n+1)$-manifold
with $\cat M=\dim M$ and $\pi_1(M)=\Zp$, but there are no closed
$2n$-manifolds with $\cat M=\dim M$ and $\pi_1(M)=\Zp$.
\end{prop}

\begin{proof}
An example of $(2n+1)$-manifold is the quotient space $S^{2n+1}/\Zp$
with respect to a free $\Zp$-action on $S^{2n+1}$. Now, given a
$2n$-manifold with $\pi_1(M)=\Zp$, consider a map $f:M \to K(\Zp,1)$
that induces an isomorphism of fundamental groups. Since
$H_{2n}(K(\Zp,1))=0$, it follows from the obstruction theory and
Poincar\'e duality that $f$ can be deformed into the $(2n-1)$-skeleton
of $K(\Zp,1)$, cf.~\cite[Section 8]{Bab1}. Hence, $M$ is not
$2n$-essential, and thus $\cat M <2n$ \cite{KR1}.
\end{proof}

%\vfill\eject

\end{document}